OpenStudy (anonymous):

Solve the given equation 5x^2 - 6x +3 =0

OpenStudy (xxdragoonxx):

x^2 - 6x/5 + 3/5 = 0 Divide both sides by 5. Subtract 3/5 from both sides. x^2 - 6x/5 = - 3/5 Add (-6/5/2)2 to both sides. x^2 - 6x/5 + 9/25 = - 6/25 Complete left hand side to a square using the identity. (x-3/5^2 = -6/25